Max    u=q1^6 q2^4+1.5Ln(q1 )+Ln(q2)

s.to

3q1+4q2=100


1
Expert's answer
2020-12-28T09:04:20-0500

According to the question it is necessary to maximize the function "u=q_{1}^6q_{2}^4+1.5ln(q_{1})+ln(q_{2})" taking into account the following restriction equation "3q_{1}+4q_{2}=100"

ANSWER

The first step is the Langrangian function (L) creation.

This function is the following:

"L(q_{1},q_{2},\lambda)=u(q_{1},q_{2})+\lambda*(3q_{1}+4q_{2}-100)"

The second step is to take the partial derivatives of the Lagrangian with respect to q1 and q2. After that using the restriction equation and the equations that the partial Langrangian derivatives are equal to zero we can find the stationary points of L.

"dL/dq_{1}=6*q_{1}^5q_{2}^4+1.5/q_{1}+3\lambda=0" (1)

"dL/dq_{2}=4q_{1}^6q_{2}^3+1/q_{2}+4\lambda=0" (2)

We can multyply the equation (2) by a q2:

"4q_{1}^6q_{2}^4+1+4\lambda*q_{2}=0" (2.1)

We can multyply the equation (1) by a q1:

"6*q_{1}^6q_{2}^4+1.5+3\lambda*q_{1}=0" (1.1)

The equation (1.1) can be rewritten as follows:

"q_{1}^6q_{2}^4=-1.5/6-3\lambda*q_{1}/6=-0.25-\lambda*q_{1}/2" (1.2)

Taking into account the (1.2) and (2.1) we can write:

"-1-2\lambda*q_{1}+1+4\lambda*q_{2}=0"

and hence "q_{1}=2q_{2}" (1.3)

Using the restriction equation we can find q1 and q2

"3*2q_{2}+4q_{2}=100"

Hence, "q_{2}=10" and "q_{1}=20". It is stationary point (that does not depend on "\lambda" ).

The third step

To prove that the values q2=10 and q1=20 are solution we need to prove that "d_{2}L" is negative in the point q2=10 and q1=20

"d^2L=d^2L\/dq_{1}^2*dq_{1}^2+2d^2L\/(dq_{1}dq_{2})*dq_{1}dq_{2}+d^2L\/dq_{2}^2*dq_{2}^2" (3.1)


"d^2L\/dq_{1}^2=30q_{1}^4q_{2}^4-1.5\/q_{1}^2=a" (3.2)

"d^2L\/dq_{2}^2=12q_{1}^6q_{2}^2-1\/q_{2}^2=b" (3.3)

"2d^2L\/(dq_{1}dq_{2})=24q_{1}^5q_{2}^3=c" (3.4)

Taking into account (3.2), (3.3) and (3.4) the equation (3.1) can be rewritten as follows:

"d^2L=a*dq_{1}^2+2b*dq_{1}dq_{2}+c*dq_{2}^2" (3.1.1)

Based on the restriction equation we can write that

"3dq_{1}=-4dq_{2}"

Hence. (3.1.1) can be rewritten as follows

"d^2L=a*dq_{1}^2-2*3\/4b*dq_{1}dq_{1}+c*dq_{1}^2*9\/16" (3.1.2)

Then,

"d^2L=(a-1.5b+c*9\/16)dq_{1}^2" (3.1.2).

Let's calculate the a, b, and c while q2=10 and q1=20.

When q2=10 and q1=20, "a=30*100*160,000-1.5\/400=480,000,000-1.5\/400"

When q2=10 and q1=20, "b=12*64,000,000*100-0.01=76,800,000,000-0.01"

When q2=10 and q1=20, "c=24*1,000*3,200,000= 76,800,000,000"


Based on the values of a, b, and c we can conclude that while q2=10 and q1=20 the "d_{2}L" is less than zero. Hence, the point q2=10 and q1=20 is point where the value of the utilitu function (denoted by u) is the highest.